# Preference Distillation for Personalized Generative Recommendation

### v1

@@
Abstract
Recently, researchers have investigated the capabilities of Large Language Models (LLMs) for generative recommender systems. Existing LLM-based recommender models are trained by adding user and item IDs to a discrete prompt template. However, the disconnect between IDs and natural language makes it difficult for the LLM to learn the relationship between users. To address this issue, we propose a PErsonAlized PrOmpt Distillation (PeaPOD) approach, to distill user preferences as personalized soft prompts. Considering the complexities of user preferences in the real world, we maintain a shared set of learnable prompts that are dynamically weighted based on the user’s interests to construct the user-personalized prompt in a compositional manner. Experimental results on three real-world datasets demonstrate the effectiveness of our PeaPOD model on sequential recommendation, top-n recommendation, and explanation generation tasks.
Preference Distillation for Personalized Generative Recommendation
Jerome Ramos  Bin Wu  Aldo Lipani University College London, London, UK {jerome.ramos.20, bin.wu.23, aldo.lipani}@ucl.ac.uk
1 Introduction
@@
Recommender systems Ge et al. (2022) have emerged as pivotal tools in guiding users through vast amounts of items and tailoring personalized experiences across various tasks, including sequential recommendation, top-n recommendation and explanation generation. Recently, recommender systems trained using large language models (LLMs), known as generative recommender systems, have unlocked the ability to perform multiple tasks with a single model. It has been shown to perform well in a zero- and few-shot setting Sanner et al. (2023) and can be further fine-tuned on a specific domain to improve performance  Ramos et al. (2024).
@@
P5  Geng et al. (2022) formulates recommendation as a text-to-text sequence task and trains the model to perform generative recommendation based on discrete prompt templates. User and item IDs serve as important identifiers for discrete prompt templates, beyond the simple semantic meaning within their word embeddings. Thus, the disconnect between IDs and natural language and the inherent rigidity of discrete prompts makes it difficult for the LLM to learn the relationship between users, limiting the effectiveness of this approach. Prompt Distillation has been proposed to distill the knowledge in discrete prompts into continuous prompt vectors which do not map to any specific words Li et al. (2023). However, the existing work distills the knowledge into global-shared prompts, which is in conflict with the fact that users’ personalized preferences greatly differ across the entire user population. Given that each user has their own unique preferences (Wu et al., 2022), it is imperative to design a method to effectively distill preferences for personalized recommendation.
One intuitive way to introduce distilled preferences is to create unique prompts per user based solely on their individual interaction history. However, there are two key issues: (1) Distilling separated prompts for each user is costly and difficult, where the number of prompts increases with the number of the user and in most cases, each user’s interaction history is too sparse to effectively personalize the recommender; and (2) This approach does not allow the model to match users with similar interests, which has been shown to increase performance for traditional recommender models in a warm-start setting Koren et al. (2009); Mnih and Salakhutdinov (2007); Cheng et al. (2016).
In this work, we propose PErsonAlized PrOmpt Distillation (PeaPOD), a more flexible and effective framework to distill user preferences as personalized soft prompts. Specifically, we construct a set of decomposed prompt components, which maintains a many-to-many mapping between user preferences and continuous prompts. This is due to the compositional nature of user preferences, meaning that users might share similar interests but maintain their distinction as well. By updating only a small set of prompt components, we can efficiently share prompts among users rather than maintaining a unique prompt per user.
Moreover, to construct the many-to-many mapping (Wu et al., 2023) between users and the distilled prompt components, we propose a compositional strategy to dynamically weight the prompts based on user preference. First, we use matrix factorization Koren et al. (2009) to generate a user embedding, which compresses a user’s interaction history into a high-dimensional space. The user embeddings are used to identify user preferences, helping to create a weighted summation of the decomposed prompt components, called the user-personalized prompt. Importantly, the construction of the user-personalized prompt are dynamic across users but only maintains a limited set of shared prompt components. From experiments across three datasets, PeaPOD is able to achieve state-of-the-art performance in three recommendation tasks. We provide all data and code used at: https://github.com/jeromeramos70/peapod.
In summary, our key contributions are:
  1. 1.
    We provide a novel architecture, namely PErsonAlized PrOmpt Distillation (PeaPOD), to distill complex user preferences into a limited set of learnable prompts.
  2. 2.
    We propose a compositional strategy, dynamically weighting the decomposed prompt components to construct the user-specific prompt based on user preference.
  3. 3.
    We conduct extensive experiments, showing the effectiveness and superiority of our proposed PeaPOD model.
2 Related Works
2.1 LLMs for Recommendation
@@
One advantage of LLMs is that they are pre-trained with a large amount of data from various domains. Thus,  Sanner et al. (2023) crowdsourced a dataset of natural language summaries of user preferences and showed that prompting a language model with these user summaries yielded competitive performance with traditional baselines in a near cold-start setting.  Ramos et al. (2024) showed that a generative recommendation system trained solely on natural language user profiles are also competitive with traditional recommender systems in a warm-start setting, with the added benefit of transparency and scrutability. In addition, the zero-shot prompting of GPT-4 has also beaten SOTA models in a conversational recommender setting He et al. (2023). Although prompting LLMs can lead to strong performance in recommender tasks, they are heavily reliant on better content knowledge rather than collaborative knowledge He et al. (2023); Wu et al. (2024). Thus, if the item metadata is not already learned in the pre-trained knowledge base or if the target dataset users interact differently than the users that the LLM was originally trained on, the performance of the recommender system will suffer.
@@
To mitigate these issues, the model can be further fine-tuned on target datasets in order to train the model based on the given user population and add or update its knowledge on the item dataset. P5 Geng et al. (2022) is the first work that trains a T5-based Raffel et al. (2020) foundational recommender model to unify various recommendation tasks in a shared framework. VIP5 Geng et al. (2023) extended P5 to a multi-modal foundational recommender system. POD Li et al. (2023) builds upon the P5 model and appends a task-specific, continuous prompt to the input to improve performance. Although these models pass in the user_id as inputs, they still do not leverage the collaborative information contained from similar users. In particular, they do not implement a user-specific component to identify and share knowledge between users with similar preferences. Previous work in continual learning for image classification has shown that an attention-based mechanism based on the image input is able to learn similarities across similar images to achieve state-of-the-art performance Smith et al. (2023). In this work, we focus on adding an attention-based mechanism to our model in order to leverage the shared collaborative knowledge between users to improve the recommendation performance.
2.2 Structural Prompt
@@
One issue with generative approaches to recommendation is how to effectively encode user-item interactions to use with LLMs. Researchers have proposed numerous methods to bridge the gap between item IDs and natural language. One option is to add the user’s name and item title to the input prompt. However, this data may be unavailable or non-unique, making it difficult to personalize recommendations. Another option is to directly insert the user and item IDs to the discrete prompt template. However, LLMs are trained on natural language text rather than IDs, which limits the effectiveness of this approach.
@@
To tackle these issues, many researchers have explored how to effectively encode item IDs with LLMs. For example, Rajput et al. (2023) proposed using RQ-VAE with item metadata to generate semantically meaningful tuples of tokens to serve as the Semantic ID for each item. GPTRec uses a novel SVD tokenization method to generate quantized item embeddings based on user-item interactions Petrov and Macdonald (2023).  Tan et al. (2024) proposed training a textual ID generator alongside the LLM-based recommender to convert item ids to meaningful text descriptions. Although these approaches have improved item IDs encoding for generative recommendations, the effective encoding of user information remains largely unexplored. In particular, existing works do not tackle the problem of how to encode both a user’s personalized preferences and the shared knowledge between similar users to an LLM. Thus, our work focuses on how to develop user-personalized prompts to improve the performance of a generative recommender system.
3 Preliminary
3.1 Recommendation Tasks
@@
Following the existing works Geng et al. (2022), we focus mainly on three important recommendation tasks: (1) Sequential Recommendation, (2) Top-n Recommendation, and (3) Explanation Generation. Let $u\in U$ denote a user in the user set and $i\in I$ denote an item in the item set across all three tasks. We aim to train a unified model to complete the following tasks:
@@
3.2 Prompt Distillation
@@
In traditional recommender systems, user and item IDs are commonly used to identify unique users and items. However, since these IDs are unique to the dataset and are not natural language text, it is necessary to perform additional fine-tuning to allow the LLM to effectively understand these IDs. The most common methodology is to train the model with discrete prompt templates. For example, in the case of explanation generation, the discrete prompt could be “Explain why user_1234 enjoys item_5678”. When using IDs with discrete prompt templates, recent research proposes adding an additional embedding space, called whole-word embeddings, so that the model can recognize which tokens belong to the appropriate ID Geng et al. (2022). In particular, each ID consists of the same whole-word embedding and all other tokens are tokenized as normal.
@@
Although discrete prompts can lead to good performance, they are inflexible and require manually crafted prompts. Thus, the performance of the model may depend strongly on how well the instructions were written. On the other hand, continuous prompts are dynamically adjusted and fine-tuned through techniques like soft prompt tuning, which optimizes prompts by learning continuous embeddings. This approach allows the model to learn regardless of the the quality of the prompt template. Li et al. (2023) introduced Prompt Distillation, which distills discrete prompts into continuous prompt vectors. The authors used prompt distillation to add task-specific continuous prompts per recommendation task globally shared across all users to improve the model’s performance. Since this method adds a shared continuous prompt per recommendation task, we will refer to these continuous prompts as task-specific prompts. Although task-specific prompts were shown to improve performance in our recommendation tasks, they cannot express personalization because each user uses the exact same prompts. Thus, it is imperative to construct a new method to distill each user’s preferences into a personalized soft prompt.
4 User-Personalized Prompt
Figure 1: Method to create the user-personalized prompt. The user embedding from matrix factorization is used to weight the decomposed prompt components. The distillled user preferences are then compositionally combined into a single soft prompt that can be passed to the model as input.
Figure 2: Full PeaPOD model architecture. PeaPOD prepends two continuous prompts to the discrete prompt template: (1) a user-personalized prompt based on the user’s preferences and (2) a global task-specific prompt shared across all users. We add the additional whole-word embedding to distinguish tokens belonging to a given ID and train an encoder-decoder model.
The goal of the user-personalized prompt is to distill user preferences into a soft prompt that is able to represent the user’s personal interests while also leveraging collaborative knowledge between similar users. We maintain a set of decomposed prompt components, which represents a many-to-many mapping between user preferences and continuous prompt vectors. This allows us to efficiently maintain a limited group of prompts and share knowledge across similar users. The decomposed prompt vectors are then dynamically weighted and combined based on the user’s interaction history to form a single continuous prompt. Using this methodology, user with similar preferences should have similar weight distributions. For example, if a decomposed prompt component contained information about action movies, then all users with an affinity to actions movies will contribute significant weight to this prompt component.
4.1 Query Function
Since we want to weight the decomposed prompt components based on each user’s preferences, we introduce a query function $q(x)$. The purpose of $q(x)$ is to obtain an embedding that effectively captures a user’s personalized interests. In our case, we choose to use a user embedding generated from matrix factorization Koren et al. (2009) as the query function. In particular, given a user-item feedback matrix $A\in\mathbb{R}^{|U|\times|I|}$, the model learns a user embedding $U\in\mathbb{R}^{|U|\times D}$ and item embedding $U\in\mathbb{R}^{|I|\times D}$, where $D$ is the dimension size and $|U|$ and $|I|$ are the number of users and items, respectively. By using a user embedding as the query function, we can effectively encode user preferences into the shared set of prompt components.
4.2 Prompt-Component Weighting
We use an attention-based architecture Vaswani et al. (2017) to attend the prompt components based on the user’s personalized preferences. Let $A\in\mathbb{R}^{M\times D}$ be a learnable attention vector, where $M$ is the number of decomposed prompt components. We calculate the element-wise multiplication between q(x) and $A$ to obtain the attended query, denoted as $\lambda$. The attended query allows the model to focus on the most important aspects of the user’s interaction history. To obtain the attended query we calculate:
$\lambda=q(x)\odot A$ (1)
where $\odot$ is the element-wise product. We then calculate the cosine similarity, denoted as $\gamma$, between $\lambda$ and $K\in\mathbb{R}^{M\times D}$, where $k_{M}\in K$ is a learnable key that has a corresponding attention vector $A_{M}$ and prompt component $p_{M}$. The cosine similarity between the attended query $\lambda$ and $K$ is used to determine the magnitude that each prompt component should receive, which we define as:
$\alpha=\gamma(\lambda,K)$ (2)
where $\alpha\in\mathbb{R}^{M\times D}$ is a weighting vector. Since $\alpha$ is updated during training, the model is able to update its representation of the user embedding rather than fixing it to a single representation. $\alpha$ can then be used to weight the prompt components based on attended query function. Thus, we can determine how much each preference should contribute to the user’s overall user-personalized prompt.
4.3 Prompt Composition
Finally, we introduce the user-personalized prompt $P$, which is calculated as a weighted summation over the set of decomposed prompt components:
$P=\sum_{M}\alpha_{M}p_{M}$ (3)
where $p\in\mathbb{R}^{L\times D\times M}$ is a decomposed prompt component and $L$ is the length of each prompt component. Each decomposed prompt component is a soft prompt that represents information that can be collaboratively shared across similar users. We base this idea on the assumption that users may share similar interests and this shared collaborative knowledge can be used to improve personalization. Our many-to-many mapping between user preferences and soft prompts are dynamically weighted using $\alpha$ based on each user’s interactions, distilling user preferences to a user-personalized prompt in a compositional manner. Notably, our model can capture the complexities of a personalized user profile while maintaining a fixed number of prompts. Since the user-personalized prompt focuses on the most important parts of the user embedding, our method is similar high dimension clustering, meaning that users with similar preferences will have similar user-personalized prompts. We show how to obtain the user-personalized prompt in Figure 1.
4.4 Orthogonal Initialization
To reduce interference in knowledge between prompt components, we can initialize $P$, $K$, and $A$ orthogonally using QR Decomposition. Namely, let $Z\in\mathbb{R}^{R\times C}$ be a real matrix with $R$ rows and $C$ columns. $Z$ can be broken into an orthogonal matrix $Q$ and upper matrix $R$. The columns of $Q$ will thus be a set of $C$ orthogonal unit vectors which we can use as the initial vectors. We can repeat this initialization process for matrices $P$, $K$, and $A$.
4.5 Training and Generative Recommendation

### v2

# PeaPOD: Personalized Prompt Distillation for Generative Recommendation

@@
Abstract.
Large Language Models (LLMs) have become a powerful foundation for generative recommender systems, offering a unified approach to diverse recommendation tasks. However, existing multi-task LLM recommendation methods often rely on discrete ID-based prompts or task-specific soft prompts, which overlook the valuable collaborative signals shared among users with similar interests. To address this limitation, this paper presents a compositional framework that integrates a user’s individual preferences with collective preferences from similar users to build personalized soft prompts. Specifically, an attention-based mechanism fuses the embeddings of users with similar interests, creating a richer representation that captures multiple facets of user preferences. This design dynamically emphasizes shared interests while maintaining each user’s distinctive preferences. Experiments on three real-world datasets demonstrate the effectiveness of the proposed approach across sequential recommendation, top-n recommendation, and explanation generation tasks, underscoring the advantages of incorporating collaborative signals through an attention-based compositional strategy.
Recommender Systems, Personalization, Generative Recommendation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
@@
Recommender systems have emerged as pivotal tools in guiding users through vast amounts of items and tailoring personalized experiences across various tasks, including sequential recommendation, top-n recommendation, and explanation generation (Ge et al., 2022). Recently, recommender systems trained using large language models (LLMs), known as generative recommender systems, have unlocked the ability to perform multiple tasks with a single model. This unified design allows LLM-based recommender systems to perform well in various downstream tasks, including in a zero-shot or few-shot setting (Geng et al., 2022, 2023).
@@
P5  (Geng et al., 2022) formulates recommendation as a text-to-text sequence task and trains the model to perform generative recommendations based on discrete prompt templates. Li et al. (2023) further proposed Prompt Distillation to address the disconnect between the IDs (including user and item) used in the discrete template and their original natural language meaning. In particular, they employ a continuous prompt to capture the information of the discrete prompt in a more flexible manner. However, existing work maintains prompts that are reused for all users. This globally-shared prompt design is in conflict with the fact that users’ personalized preferences greatly differ across the entire user population (Wu et al., 2022). In the real world, users might have similar or distinct preferences behind their interactions.
One intuitive solution for personalizing generative recommender systems is to create a personalized prompt for each user based solely on their individual interactions.  Tan et al. (2024b) have shown that fine-tuning a unique parameter-efficient fine-tuning (PEFT) module on each user’s interaction history can effectively personalize LLMs for personalized text generation and classification. However, this method is is reliant on the user having extensive historical data. Thus, fine-tuning a unique prompt per user is ineffective when the user’s interaction history is too sparse to enable effective personalization. Additionally, prior research in recommender systems has shown that sharing collaborative information among users with similar interests can improve performance in warm-start settings (Koren et al., 2009; Mnih and Salakhutdinov, 2007; Cheng et al., 2016).
Figure 1. Comparison of soft prompt architectures for multi-task generative recommendation. Unlike previous methods, PeaPOD utilizes the query-key-value mechanism from attention to generate a collaborative user prompt that captures group-level preferences based on the target user’s individual preferences.
In this work, we propose PErsonAlized PrOmpt Distillation (PeaPOD), a flexible framework that fuses individual and group-level signals into a unified collaborative user prompt for recommendation. To construct this system, we first utilize standard collaborative filtering techniques to generate initial user embeddings. These embeddings are then refined using a compositional, attention-based strategy that captures a contextualized representation of user-to-user relationships. Specifically, the user embedding of target user $u$ serves as the query vector in a multi-head attention module, while the embeddings of the top-n most similar users, identified via cosine similarity, act as both key and value vectors. By dynamically weighting shared preferences among similar users, our model generates a collaborative user soft prompt that effectively captures and aligns with the target user’s preferences. Moreover, we implement multiple attention heads to allow the model to capture different dimensions of similarity, resulting in a richer representation of the target user. As a result, PeaPOD provides a personalized soft prompt that can be used effectively across multiple generative recommendation tasks. Across three Amazon datasets, PeaPOD is able to achieve state-of-the-art performance on three recommendation tasks. We provide all data and code used at: https://anonymous.4open.science/r/peapod-3FF7/.
Our key contributions are summarized below:
  1. (1)
    We propose PErsonAlized PrOmpt Distillation (PeaPOD), a novel compositional strategy that distills collaborative user preferences into a personalized soft prompt for LLM-based recommendation.
  2. (2)
    We leverage the query-key-value mechanism from multi-head attention to capture diverse user-to-user relationships and construct a contextualized user representation that integrates both individual- and group-level information.
  3. (3)
    We conduct extensive experiments on sequential recommendation, top-n recommendation, and explanation generation tasks. Comprehensive ablation studies and results further validate the effectiveness and superiority of our proposed PeaPOD framework.
2. Related Works
2.1. LLMs for Recommendation
@@
One advantage of LLMs is that they are pre-trained with a large amount of data from various domains. Thus,  Sanner et al. (2023) crowdsourced a dataset of natural language summaries of user preferences and showed that prompting a language model with these user summaries yielded competitive performance with traditional baselines in a near cold-start setting.  Ramos et al. (2024) showed that a generative recommendation system trained solely on natural language user profiles are also competitive with traditional recommender systems in a warm-start setting, with the added benefit of transparency and scrutability. In addition, the zero-shot prompting of GPT-4 has also beaten SOTA models in a conversational recommender setting (He et al., 2023). Although prompting LLMs can lead to strong performance in recommender tasks, they are reliant on better content knowledge rather than collaborative knowledge (He et al., 2023; Wu et al., 2024). Thus, if the item metadata is not already learned in the pre-trained knowledge base or if the target dataset users interact differently than the users that the LLM was originally trained on, the performance of the recommender system will suffer.
@@
To mitigate these issues, the model can be further fine-tuned on target datasets in order to train the model based on the given user population and add or update its knowledge on the item dataset. P5 (Geng et al., 2022) is the first work that trains a T5-based (Raffel et al., 2020) foundational recommender model to unify various recommendation tasks in a shared framework. VIP5 (Geng et al., 2023) extends P5 to a multi-modal foundational recommender system. POD (Li et al., 2023) builds upon the P5 model and appends a task-specific, continuous prompt to the input to improve performance. Although these models pass in the user ID as inputs, they still do not leverage the collaborative information contained from similar users. Zhu et al. (2024) expands the tokenizer vocabulary to add user/item IDs as tokens and learn collaborative information, but this technique is infeasible as the number of users/items grows. In this work, we focus on developing a foundational generative recommender model that can efficiently encode individual and collaborative preferences for multiple downstream tasks.
2.2. Structural Prompt
@@
One issue with generative approaches to recommendation is how to effectively encode user-item interactions to use with LLMs. Researchers have proposed numerous methods to bridge the gap between IDs and natural language. One option is to add the user’s name and item title to the input prompt. However, this data may be unavailable or non-unique, making it difficult to personalize recommendations. Another option is to directly insert the user and item IDs to the discrete prompt template (Geng et al., 2022). However, LLMs are trained on natural language text rather than IDs, which limits the effectiveness of this approach (Li et al., 2023).
@@
To tackle these issues, many researchers have explored how to effectively encode item IDs with LLMs.  Rajput et al. (2023) proposed using RQ-VAE with item metadata to generate semantically meaningful tuples of tokens to serve as the semantic ID for each item. GPTRec uses a novel SVD tokenization method to generate quantized item embeddings based on user-item interactions (Petrov and Macdonald, 2023).  Tan et al. (2024a) proposed training a textual ID generator alongside the LLM-based recommender to convert item IDs to meaningful text descriptions. While these methods are effective, they do not address the need for encoding user preferences as input to the model.
To incorporate user preferences into the LLM-based recommender, Zhang et al. (2024) proposes a mapping technique to project user embeddings learned from collaborative filtering into a soft prompt. In this work, we focus on developing an attention-based compositional method to better model user-to-user relationships. Previous work in continual learning for image classification has shown that an attention-based mechanism can be used to effectively share information between similar classes of images to improve performance (Smith et al., 2023). In the context of generative recommendation, multi-head attention (Vaswani et al., 2017) is used to capture both individual and collaborative preferences across diverse feature spaces.
3. Preliminary
3.1. Recommendation Tasks
@@
Following existing works (Geng et al., 2022; Li et al., 2023), we focus on three important recommendation tasks: (1) Sequential Recommendation, (2) Top-n Recommendation, and (3) Explanation Generation. Let $u\in U$ denote a user in the user set and $i\in I$ denote an item in the item set across all three tasks. We aim to train a unified model to complete the following tasks:
@@
Table 1. Notation Table
Symbol Description
$u$ Target user
$i$ Item
$E_{u,i}$ Explanation why $u$ likes/dislikes $i$
$\boldsymbol{F}$ User-Item feedback matrix
$d_{u}$ Dimension of user embedding
$S_{u}$ Top-n most similar users to $u$
$d_{m}$ Dimension of model
$\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ Query, key, and value matrices
$\mathbf{W}_{h}^{Q}$, $\mathbf{W}_{h}^{K}$, $\mathbf{W}_{h}^{V}$ Linear projection weights for head $h$
$\mathbf{z}$ User’s summary vector
$\mathbf{p}$ Collaborative user prompt
$d_{p}$ Dimension of collaborative user prompt
$\boldsymbol{head}_{h}$ Output of attention for head $h$
$d_{k}$ Dimension of attention head
$\mathbf{W}_{l}$ Linear layer weight for feedforward network
$\mathbf{b}$ Bias vector
$H$ Number of attention heads
3.2. Prompt Distillation
@@
In traditional recommender systems, user and item IDs are commonly used to identify unique users and items. However, since these IDs are unique to the dataset and are not natural language text, it is necessary to perform additional fine-tuning to allow the LLM to effectively understand these IDs. The most common methodology is to train the model with discrete prompt templates. For example, in the case of explanation generation, the discrete prompt could be “Explain why user_1234 enjoys item_5678”. When using IDs with discrete prompt templates, recent research proposes adding an additional embedding space, called whole-word embeddings, so that the model can recognize which tokens belong to the appropriate ID (Geng et al., 2022). That is, each token in the ID utilizes the same whole-word embedding and all other tokens are tokenized as normal.
@@
Although discrete prompts can lead to good performance, they are inflexible and require manually crafted prompts. Thus, the performance of the model may depend strongly on how well the instructions were written. On the other hand, continuous prompts are dynamically adjusted and fine-tuned through techniques like soft prompt tuning, which optimizes prompts by learning continuous embeddings. This approach allows the model to learn regardless of the quality of the prompt template. Li et al. (2023) introduced Prompt Distillation, which distills discrete prompts into continuous prompt vectors. The authors used prompt distillation to incorporate task-specific, continuous prompts for each recommendation task. These prompts are shared globally across all users to improve the model’s performance. Since this method adds a shared continuous prompt per recommendation task, we will refer to these continuous prompts as task-specific prompts. Although task-specific prompts were shown to improve performance in our recommendation tasks, they cannot express personalization because each user utilizes the same set of prompts. Thus, it is imperative to develop a new method to distill each user’s preferences into a personalized soft prompt that caters to each user’s individual needs, while incorporating relevant preferences from similar users.
4. Collaborative User Prompt
Figure 2. Full PeaPOD model architecture. We first initialize a set of user embeddings using probabilistic matrix factorization. We then apply multi-head attention, where the user embedding of target user $u$ is passed to the query vector and the embeddings of the top-n most similar users $S_{u}$ are passed to the key and value vectors. The final output is the collaborative user prompt $\boldsymbol{p}_{mha}$, which dynamically captures collaborative knowledge from $S_{u}$ based on $u$’s personalized preferences.
The goal of the collaborative user prompt is to distill user preferences into a soft prompt to represent the user’s personal interests while also leveraging collaborative knowledge from users with similar preferences. To accomplish this, we implement multi-head attention (Vaswani et al., 2017) to create a contextualized user representation that combines individual and group-level information. Attention dynamically weights the influence of each similar user, producing a richer and more robust representation. Furthermore, the utilization of multiple heads allows the model to capture different aspects of similarity, resulting in a more comprehensive and diverse understanding of the target user. By applying multi-head attention, we are able to generate a personalized soft prompt that can be effectively used and shared by the LLM across multiple recommender tasks.
4.1. User Embedding Representation
To generate the collaborative user prompt, we must first select an effective input to the multi-head-attention module. In our case, we choose to utilize a user embedding generated from probabilistic matrix factorization (Mnih and Salakhutdinov, 2007) to generate embeddings for each user. In particular, given a user-item feedback matrix $\boldsymbol{F}\in\mathbb{R}^{|\boldsymbol{U}|\times|\boldsymbol{I}|}$, the model learns a user matrix $\boldsymbol{U}\in\mathbb{R}^{|\boldsymbol{U}|\times d_{u}}$ and item matrix $\boldsymbol{I}\in\mathbb{R}^{|\boldsymbol{I}|\times d_{u}}$, where $d_{u}$ is the selected dimension size of the user embedding and $|\boldsymbol{U}|$ and $|\boldsymbol{I}|$ are the number of users and items, respectively. $\boldsymbol{u}\in\boldsymbol{U}$ represents a user embedding that encodes a user’s individual preferences. By construction, users with similar preferences end up closer in this embedding space (e.g. higher cosine similarity). These pretrained user embeddings serve as a strong initialization for personalization because they already capture collaborative signals.
4.2. Attending User Embeddings
In our framework, attention is utilized to generate a collaborative user prompt $\boldsymbol{p}$ by leveraging the embeddings of user $u$ (e.g. $\boldsymbol{u}$) and $\boldsymbol{S}_{u}\in\mathbb{R}^{n\times d_{u}}$, which is defined as the top-n most similar user embeddings to $\boldsymbol{u}$ by cosine similarity. We attend over these $n$ neighboring user embeddings to capture their collective preferences, conditioned on $u$. We first project $\boldsymbol{u}$ and each neighbor’s embedding $\boldsymbol{s}_{i}\in\boldsymbol{S}_{u}$ into the query $\boldsymbol{Q}$, key $\boldsymbol{K}$, and value $\boldsymbol{V}$ subspaces.
(1) $\boldsymbol{Q}$ $=\boldsymbol{W}_{Q}u+\boldsymbol{b}_{Q}\in\mathbb{R}^{d_{m}}$
(2) $\boldsymbol{K_{i}}$ $=\boldsymbol{W}_{K}s_{i}+\boldsymbol{b}_{K}\in\mathbb{R}^{d_{m}}$
(3) $\boldsymbol{V}_{i}$ $=\boldsymbol{W}_{V}s_{i}+\boldsymbol{b}_{V}\in\mathbb{R}^{d_{m}}$
where $d_{m}$ is the dimensionality of the base LLM. Aggregating similar users, we define the key and value spaces as:
(4) $\mathbf{K}=\begin{bmatrix}\mathbf{K}_{1}\\ \mathbf{K}_{2}\\ \vdots\\ \mathbf{K}_{n}\end{bmatrix}\in\mathbb{R}^{n\times d_{m}},\quad\mathbf{V}=\begin{bmatrix}\mathbf{V}_{1}\\ \mathbf{V}_{2}\\ \vdots\\ \mathbf{V}_{n}\end{bmatrix}\in\mathbb{R}^{n\times d_{m}}$
$\boldsymbol{Q}$ is the representation of the target user $u$ that we condition group-level information on. $\boldsymbol{K}$ encodes $\boldsymbol{S}_{u}$ in the key space, enabling the model to learn how relevant each similar user is to $u$. $\boldsymbol{V}$ carries the content from $\boldsymbol{S}_{u}$ to be aggregated by attention. We use scaled dot-product attention to compute an attention vector and apply it to $\boldsymbol{V}$:
(5) $\boldsymbol{z}=\text{Attention}(\boldsymbol{Q,K,V})=\text{softmax}\left(\frac{\boldsymbol{QK}^{\top}}{\sqrt{d_{m}}}\right)\boldsymbol{V}\in\mathbb{R}^{d_{m}},$
This results in an attention-weighted summary vector $\boldsymbol{z}$. In the context of our selected inputs, This allows us to dynamically aggregate the most important information from similar users with regards to $u$, further enhancing the collaborative information learned from the initialized user embeddings. Finally, we apply a linear layer to $\boldsymbol{z}$ to obtain the collaborative user prompt $\boldsymbol{p}$.
(6) $\boldsymbol{p}$ $=\boldsymbol{W}_{l}\boldsymbol{z}+\boldsymbol{b}_{l}\in\mathbb{R}^{d_{p}}$
where $d_{p}$ is the selected prompt length for the collaborative user prompt. By attending to the top-n most similar users, we dynamically incorporate relevant collaborative signals into $\boldsymbol{p}$ based on $u$. This compositional strategy combines group-level information, conditioned on $u$’s preferences, to distill collaborative preferences into a personalized soft prompt.
4.3. Multi-Head Attention
The collaborative user prompt described in Section 4.2 can be extended from a single-head implementation to a multi-head attention (Vaswani et al., 2017). Multi-head attention enables the model to attend to different subspaces of the embeddings simultaneously, allowing the model to capture richer relationships and patterns across the top-$n$ similar users.
Formally, instead of using a single set of linear projections for $\boldsymbol{Q}$, $\boldsymbol{K}$, and $\boldsymbol{V}$, we learn multiple sets of projections, one for each head $h\in\{1,\dots,H\}$:
(7) $\textbf{head}_{h}$ $=\text{Attention}\bigl{(}\boldsymbol{QW}_{h}^{Q},\;\boldsymbol{KW}_{h}^{K},\;\boldsymbol{VW}_{h}^{V}\bigr{)}\in\mathbb{R}^{d_{k}},$
where $\mathbb{R}^{d_{k}}=\mathbb{R}^{d_{m}/H}$ is the dimension of each attention head. After concatenating all heads, the final dimension remains $d_{m}$. The outputs of each head are then concatenated and passed through a linear layer:
(8) $\boldsymbol{z}_{mha}=\text{Concat}\bigl{(}\textbf{head}_{1},\dots,\textbf{head}_{H}\bigr{)}\in\mathbb{R}^{d_{m}}\$
(9) $\boldsymbol{p}_{mha}=\boldsymbol{W}_{l}\boldsymbol{z}_{mha}+\boldsymbol{b}_{l}\in\mathbb{R}^{d_{p}}$
In the context of the collaborative user prompt, multi-head attention is able to capture multiple, diverse user-to-user relationships between $u$ and $S_{u}$. Specifically, each head can learn to focus on different aspects of the user embeddings. For example, one head might focus on users who share highly overlapping item preferences, while another might capture niche or temporal similarities. By combining multiple heads, we integrate these different viewpoints into a more comprehensive collaborative user representation. Because multi-head attention allows us to capture more nuanced similarities among neighbors, it potentially leads to better personalization than single-head attention.
4.4. Training and Generative Recommendation